OpenStudy (anonymous):

Algebra 1 medal will be rewarded!

OpenStudy (anonymous):

using the formula y = mx+b where m is your slope and b is your y intercept, we can see that -4 is the y intercept and your slope is -1/3 

OpenStudy (anonymous):

so the answer is the 3rd graph you posted

jimthompson5910 (jim_thompson5910):

start at the point (0, -4), which is the y-intercept then go up 1 unit, then to the right 3 units (since slope = rise/run = 1/3) to get to the next point (3, -3)

jimthompson5910 (jim_thompson5910):

plot those two points, then draw a line through them to complete the graph

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

yw